# The Unseen Species Problem Revisited

### v1

@@ 4.1.1 Algorithmically Constructing Linear Estimators
Since the GTE and SGTEs are linear, this in particular means that $\hat{S}_{t,T}^{(H^{*})}$ achieves (at least) their rates and hence obtains the optimal rate for $r\leq 1$ and is minimax pseudo-optimal for $r>1$.
4.1.2 Implementation
Of course on a computer we can only minimize over a finite dimensional space and so for the purposes of producing the figures in the present paper we restrict ourselves to $H$ such that $H_{k}=0$ for all $k$ larger than $10$. To minimize $G_{H}$ we approximate the suprema using softmax and use a gradient based method (L-BFGS-B) over a gradually refining grid. Code for the solver as well as the code used to generate all the figures in the paper, is available on our GitHub 333github.com/EdwardEriksson/Unseen-Species. We will sometimes abuse terminology/notation and speak of the finite-dimensional approximation as if it was $\hat{S}_{t,T}^{(H^{*})}$.
@@ 5.2 Generalized Setting
Because our suggested estimator for alpha is invariant under the map $\boldsymbol{\phi}\rightarrow k\boldsymbol{\phi}$ the induced estimator $\hat{S}_{t,T}$ is $\phi$-homogeneous while that of [17] is not.
@@
As described on page 3285 of [3], the authors of [12] show that the bounded arity assumption in Lemma 45 can be relaxed to $\mu$ being such that there exists a size-biased coupling of $S_{t}$, which we denote by $S_{t}^{(b)}$, so that there exists a $p\in(0,1]$ such that for all $x\in\mathbb{Z}_{\geq 0}$,
$P(S_{t}^{(b)}-S_{t}\leq B|S_{t}^{(b)}\geq x)$ $\geq p,$
$P(S_{t}^{(b)}-S_{t}\leq B|S_{t}^{(b)}\leq x)$ $\geq p.$
@@
We thank Alexander Fuchs-Kreiss for adapting the techniques of [17] to give a purely asymptotic argument that the above can in fact be strengthened to $r(t)\lesssim\exp(t^{\frac{\alpha}{2}})$ and also for many valuable conversations regarding the present work. This is a factor $\log(t)$ improvement on Theorem 3 of [17] but under a slightly stronger assumption on the power-law approximation. However, if we make only the weaker assumption we end up with their rate, so our result is strictly better.
@@ 6.1 Task
For $r>1$ we compare the performance of a variety of estimators. Our prescribed task is modeled after the use case in [1]. Following [15], [23] and [20] we consider a person reading Hamlet. They know the total number of words in the play and aim to predict the total number of distinct words from reading an initial portion of the piece. Setting $\hat{S}_{T}:=S_{t}+\hat{S}_{t,T}$ we see that this is an instantiation of the unseen species problem. We study absolute relative error, namely $\frac{|\hat{S}_{T}-S_{T}|}{S_{T}}$. When it makes sense to do so we give both the performance when the data is in the true temporal order and an average over $100$ shuffles to reduce noise. The shuffling is justified in so far that we believe in exchangeability. We prefer this to sampling i.i.d. from the empirical distribution as that underestimates the number of rare species and artificially introduces exchangeability.
@@ 6.3 Benchmarks
We compare $\hat{S}_{t,T}^{(H^{*})}$, ($H^{*}$) and our suggested distant future estimator (Ratio-$\alpha$) to several competitors. A naive benchmark is obtained from the null estimator (Trivial). We include the binomial smoothing SGTE with the suggested parameters (SGT) and the power law estimator of [17] (MLE-$\alpha$). We also compare against the method of [14] (Pade-GT[2,3][2,3]) developed for applications in genome coverage. We also implement the support size estimator of [23] (Chebyshev). In their notation we set $k$ to be the corpus length and $c_{0}=0.45,c_{1}=0.5$ as their experimentally chosen values.
@@ 6.4 Results
See the result in Figure 2. We see that for the messages data the plots for the true order and the permutation average look totally different. This strongly suggests that exchangeability was severely violated and that the model is therefore not a good fit. One sees that the excellent performance of our estimators on the permutation average does not translate into good predictions under the temporal ordering. This serves as an important reminder that exchangeability is a non-trivial modelling assumption and that applications may require specialized models.
We note that since the unseen species problem resembles support estimation only for large $r$ it is natural to expect the Chebyshev-polynomial-based estimator of [23] to perform poorly for small $r$, possibly explaining the U-shaped error curve. It appears that the Padé-based estimator is not competitive with the other estimators.
We find that our distant future estimator fairly consistently outperforms that of [17] and that in as previously discussed, the worst-case improvements of $\hat{S}_{t,T}^{(H^{*})}$ over the SGT translate to improvements when new discoveries are common, but not when they are rare. Neither the linear nor the power law estimators are consistently better than the others across data sets. No doubt in part due to variations in how good the power law approximation is.

### v2

@@ 4.1.1 Algorithmically Constructing Linear Estimators
Since the GTE and SGTE are linear this in particular means that $\hat{S}_{t,T}^{(H^{*})}$ obtains the optimal rate for $r\leq 1$ and is (at least) minimax pseudo-optimal for $r>1$. Since linearity is equivalent to having the SEP, this means that $\hat{S}_{t,T}^{(H^{*})}$ is optimal up to an explicit multiplicative constant in a pre-asymptotic sense in the class of estimators which enjoy the SEP.
@@ 4.1.2 Comparing Worst Case Guarantees
A solver to compute an approximation of $H^{*}$, as well as the code to generate all figures in the paper is available on our GitHub 555github.com/EdwardEriksson/Unseen-Species. We will sometimes abuse terminology/notation and speak of the finite-dimensional approximation as if it was $\hat{S}_{t,T}^{(H^{*})}$.
@@ 5.2 Generalized Setting
Because our suggested estimator for alpha is invariant under the map $\boldsymbol{\phi}\rightarrow k\boldsymbol{\phi}$ the induced estimator $\hat{S}_{t,T}$ is representation invariant while that of [17] is not.
@@
As described on page 3285 of [3], the authors of [12] show that the bounded arity assumption in Lemma 45 can be relaxed to $\mu$ being such that there exists a size-biased coupling of $S_{t}$, which we denote by $S_{t}^{[b]}$, so that there exists a $p\in(0,1]$ such that for all $x\in\mathbb{Z}_{\geq 0}$,
$P(S_{t}^{[b]}-S_{t}\leq B|S_{t}^{[b]}\geq x)$ $\geq p,$
$P(S_{t}^{[b]}-S_{t}\leq B|S_{t}^{[b]}\leq x)$ $\geq p.$
@@
Remark 51.
Under the power law approximation $\mathbb{E}[S_{t}]\approx c\Gamma(1-\alpha)t^{\alpha}$, so if one assumed $f$ to be negligible all quantities in Theorem 50 have natural estimators and so one can use the result to construct conservative confidence intervals. Unfortunately we find that even requesting intervals with a small nominal coverage the empirical coverage ends up being essentially $100\%$, so while the result can be used to, in a principled way, give an interval which is very likely to contain the truth, it falls short of being able to produce anything calibrated.
@@
We thank Alexander Fuchs-Kreiss for adapting the techniques of [17] to give a purely asymptotic argument that the above can in fact be strengthened to $r(t)\lesssim\exp(t^{\frac{\alpha}{2}})$. This is a factor $\log(t)$ improvement on Theorem 3 of [17] but under a slightly stronger assumption on the power-law approximation. However, if we make only the weaker assumption we end up with their rate, so our result is strictly better.
We also thank him for his careful reading of earlier drafts, for scrutinizing several arguments, and for contributing helpful elaborations on a number of technical details.
@@ 6.1 Task
For $r>1$ we compare the performance of a variety of estimators. Our prescribed task is modeled after the use case in [1]. Following [15], [24] and [21] we consider a person reading Hamlet. They know the total number of words in the play and aim to predict the total number of distinct words from reading an initial portion of the piece. Setting $\hat{S}_{T}:=S_{t}+\hat{S}_{t,T}$ we see that this is an instantiation of the unseen species problem. In order to make the plots as comparable as possible we study absolute relative error, namely $\frac{|\hat{S}_{T}-S_{T}|}{S_{T}}$. When it makes sense to do so we give both the performance when the data is in the true temporal order and an average over $100$ shuffles to reduce noise. The shuffling is justified in so far that we believe exchangeability holds. We prefer this to sampling i.i.d. from the empirical distribution as that underestimates the number of rare species and artificially introduces exchangeability.
@@ 6.3 Benchmarks
We compare $\hat{S}_{t,T}^{(H^{*})}$, ($H^{*}$) and our suggested distant future estimator (Ratio-$\alpha$) to several competitors. A naive benchmark is obtained from the null estimator (Trivial). We include the binomial smoothing SGTE with the suggested parameters (SGT), which is the SGTE with the best rate guarantee, and the power law estimator of [17] (MLE-$\alpha$). We also compare against the method of [14] (Pade-GT[2,3][2,3]) developed for applications in genome coverage. We also implement the support size estimator of [24] (Chebyshev), treating the estimated support size as an estimate for $S_{T}$. We do this to investigate the difference between the unseen species problem and support size estimation. In their notation we set $k$ to be the corpus length and $c_{0}=0.45,c_{1}=0.5$ as their experimentally chosen values.
@@ 6.4 Results
See the result in Figure 2. We see that for the messages data the plots for the true order and the permutation average look totally different. This strongly suggests that exchangeability was severely violated and that the model is therefore not a good fit. One sees that the excellent performance of our estimators on the permutation average does not translate into good predictions under the temporal ordering. This serves as an important reminder that exchangeability is a non-trivial modelling assumption and that applications may require specialized models. One way to test the suitability of the model might be to artificially use an initial portion of the data as the first expedition and the second part of the data the return expedition.
We note that as expected Chebyshev-polynomial-based estimator of [24] performs poorly for small $r$ as for such $r$ the unseen species problem is very different from support size estimation.
We find that our distant future estimator fairly consistently outperforms that of [17] and that in as previously discussed, the worst-case improvements of $\hat{S}_{t,T}^{(H^{*})}$ over the SGT translate to improvements when new discoveries are common, but not when they are rare. Neither the linear nor the power law estimators are consistently better than the others across data sets. Also, despite theoretically having drastically different prediction horizons it does not appear that the power-law estimators are dominant for large $r$. No doubt the extent to which the power-law approximation is accurate is an important factor in prediction performance. See appendix F of [17] for a diagnostic tool to assess such fit.